2,5,8,.........100 terms
3,5,7...........100 terms
 
Find the sum of all the common terms of both the AP’s?

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
Hello student,
for 2,5,8,............100 terms
for 3,5,7.............100 terms
as we can see that second series hasall odd termstill 3+(99)2=201
and 1stseries has last term as 2+(99)3=299
as we can see from aboveevery alternate termof 1stseries is present in 2ndalso.(as they are odd)
So the sum ofcommon terms of Ap’s=5+11+17+.......+201
=n/2(a+l)
=34/2(5+201)
=17(206)
=3502
Thanks and regards
Shaik Aasif
askIITians faculty